OpenStudy (anonymous):

Which could be the next step in simplifying this expression? 13 + 19 + 7 A. 20 + 19 (use the commutative property) B. 32 + 20 (use the distributive property) C. 13 + 36 (use the associative property) D. 36 + 13 (use the commutative property)

OpenStudy (anonymous):

@xapproachesinfinity

OpenStudy (anonymous):

Is it A?

OpenStudy (xapproachesinfinity):

addition you don't worry how do you do it because it is commutative A is good

OpenStudy (anonymous):

13 + 19 + 7 = (13 + 19) + 7 = 32 + 7 = 39

OpenStudy (anonymous):

Thanks :)